### v2

# Lifetime measurement of the muonic atoms of enriched Si isotopes

Abstract
- Background
A muonic atom, composed of a negative muon and an atomic nucleus, undergoes two primary decay processes: muon decay and muon nuclear capture. The branching ratio between these two processes can be determined from the measured lifetime of the muonic atom. While past researches have examined the general trend of muon capture rates across different elements, experimental and theoretical investigations into the isotope dependence of the lifetime remain limited.
- Purpose
The present study aims to measure the lifetimes of the muonic atom of isotopically enriched silicon isotopes.
- Methods
The experiment was conducted at the muon facility in the Material and Life Science Facility (MLF), J-PARC. A muon beam was stopped in various target materials, including isotopically enriched 28,29,30Si. The lifetimes of the muonic atoms were measured by detecting decay electrons using a SR spectrometer. The decay spectra were analyzed by fitting with multi-exponential functions to account for contributions from the target nucleus and surrounding materials.
- Results
For the first time, the lifetimes of the muonic atom of isotopically enriched 28,29,30Si were measured. Additionally, seven other targets were studied to validate the experimental method and analysis procedure. The results were compared with theoretical models such as Primakoff, Goulard-Primakoff, and the recently-developed microscopic and evaporation model (MEM).
- Conclusions
The Primakoff and Goulard-Primakoff formulas, while reproducing certain aspects of the isotope dependence, require further refinement. The comparison with MEM calculations constrains the axial vector () and induced pseudoscalar () coupling constants. The present experiment establishes a method for measuring the lifetimes of the muonic atom and will contribute to future systematic investigations.
I Introduction
A muonic atom consists of a negative muon and an atomic nucleus. The muon at the lowest atomic state ( state) undergoes either muon decay or muon nuclear capture. The muon decay, also known as a decay-in-orbit or Michel decay, is a conversion of the muon to an electron, an electron antineutrino, and a muon neutrino, expressed as
Muon nuclear capture [1], also known as ordinary muon capture or muon nuclear absorption, is a conversion of the muon and a proton inside the nucleus to a neutron and a muon neutrino, the elemental process of which is expressed as
The capture rate of the muon capture reaction varies depending on the nucleus forming the muonic atom. In general, the capture rate increases for heavier nuclei due to an increased overlap of the muon and nuclear wave functions. The muon capture rates () are deduced by measurements of a mean lifetime of the muonic atom () as
where is the lifetime of the positive muon, viz. 2.1969811(22) s [2], and is a Huff factor [3], which takes into account a reduction in the muon decay rate due to the atomic binding energy. A comprehensive study of the capture rates was conducted, with most results summarized in Ref. [4]. However, these measurements focused primarily on natural abundance targets, and data on the capture rates for isotopically enriched targets are still limited. A significant isotope dependence of the muon nuclear capture rate was pointed out by some measurements in the past [5, 6, 7, 8, 9, 10, 11]. In general, the capture rate decreases with an increase in the mass number of isotopes, because of the Pauli exclusion principle on the muon capture process [12]. As the neutron number increases, more neutron orbitals are filled, making it more difficult for a proton to transform into a neutron at higher levels by muon capture.
Several theoretical calculations have been conducted for the muon nuclear capture rate. The Primakoff [12] and Goulard-Primakoff [13] formulas have frequently been used to estimate the capture rate. Although these formulas provide a good overall approximation, they are less accurate in predicting the isotope dependence of the capture rate, particularly due to the lack of experimental data for isotopically enriched targets [14]. Recently, a microscopic and evaporation model (MEM) was developed to describe the comprehensive muon nuclear capture process, including muon capture, direct and pre-equilibrium processes, and particle evaporations from a compound nucleus [15]. The calculation reveals a strong dependence of the capture rate on the axial vector and induced pseudoscalar coupling constants ( and , respectively). Due to the significant dependence on these constants, the capture rate calculated by MEM also has considerable uncertainty on the capture rate. Several other theoretical calculations have also been performed [16, 17, 18, 19]. However, these theories primarily focused on the global systematics of muon capture and did not discuss the detailed isotope dependence.
The present experiment was motivated by our previous measurement of production branching ratios of residual nuclei by muon capture on enriched Si isotopes [20]. As the branching ratio of each residual nucleus was given as the number per muon nuclear capture, the capture probability () is necessary to obtain the absolute branching ratio. The isotope dependence of for 28,29,30Si has not been known to date, and the Primakoff and Goulard-Primakoff formulas do not accurately predict isotope dependence of [5, 6, 7, 8, 9, 4]. Therefore, in the present experiment, we have measured , or equivalently, the mean lifetimes for 28,29,30Si.
II Experiment
The experiment was performed at the D1 area at the muon science facility (MUSE) in the Materials and Life Science Facility (MLF) in J-PARC [21]. A proton beam from the Rapid Cycling Synchrotron (RCS) with an energy at 3 GeV irradiated a rotating graphite target to produce pions. The primary beam had a double-pulse structure with a 25 Hz repetition and a power of approximately 900 kW. The negative muon beam, a decay product of negative pions, was transported through the D-beamline. The beam momentum was selected to 32 to maximize the number of muons stopped at the target and minimize that at the air and target case. The beam was then delivered to the D1 area, where the experimental apparatus was installed. The muon beam intensity was approximately sec-1.
Figure 1 shows a schematic view of the experimental setup at the D1 area. The muon beam passed through a lead collimator with a hole size of 40 mm in diameter and irradiated a target. The targets were located approximately 65 mm downstream of the exit of the lead collimator. A total of ten targets was measured in the present experiment, as listed in Table 1. The main goal of the present study is to measure the lifetimes of the muonic atom of three isotopically enriched Si targets of 28,29,30Si. The enriched targets had been used in our previous study, and details of the target form, size, weight, chemical purity, and target enrichment are given in Table 1. The targets were in metal chippings form encapsulated in an acrylic case, as schematically illustrated in the inset of Fig. 1. The effective thickness of the targets was approximately 1.0 g/cm2. The primary source of uncertainties in the lifetime measurement in the present setup was the muon decay electrons from the light elements of carbon (C), nitrogen (N), and oxygen (O), as discussed in the next section. The silver plates were used to avoid stopping the muon beam in the acrylic case. A silver mask with a hole size of 12 mm in diameter and a thickness of 0.5 mm was placed on the upstream side of the case, and a silver backing plate with a thickness of 0.5 mm was installed on the downstream side of the samples. To evaluate the effect of the chipping form of the target in the acrylic case, natural abundance silicon (natSi) samples of a self-supported metal plate and a powder form in the same acrylic case were also measured. In addition to the Si targets, the lifetimes of the muonic atoms of natMg and 27Al, which have a similar lifetime to silicon isotopes, were measured to evaluate the reliability of the measurement. For 27Al, both a self-supported plate and a plate in the acrylic case were measured to evaluate the effect of the target case. In a natAg sample measurement, the lifetimes of background components of the measurements were evaluated. The 55Mn target was measured to evaluate the reliability of lifetime measurement in the range between the lifetime of silicon isotopes and silver. Those targets had an energy loss equivalent thickness of the enriched Si targets and were used to verify the validity of the experimental and analysis procedures.
The lifetime of the muonic atom was measured by detecting the decay electrons using a SR spectrometer based on Kalliope detectors [22]. The spectrometer consisted of a total of 40 units of the Kalliope detectors. Each unit of the Kalliope detectors was comprised of 16 pairs of plastic scintillators. A coincidence of each pair defined an electron event of the muon decay. A zero-field correction was applied to avoid the muon spin relaxation and rotation affecting the measured lifetime. The residual magnetic field at the center of the spectrometer was less than 10 mG. The data acquisition system was triggered by the reference timing of the synchrotron. As the Kalliope detector system can obtain signals until 1 Mcps without saturation [23], the dead time was negligible in the present experiment with a negative muon beam. The timings of the electron event were stored from to 57 s with respect to the timing of the beam arrival [24]. The measurement times for each target are listed in Table 1.
III Analysis
The lifetimes of the muonic atom for each target were deduced by fitting the decay spectra. These spectra were produced by summing up the timings from all 40 Kalliope detector units. The timing spectra are shown in Figs. 2 and 3.
Backgrounds in the spectra were evaluated using an Ag metal plate target in the acrylic case, as shown in Fig. 2. Due to the double pulse structure of the beam, two peaks were observed in the timing spectra, along with two peaks at an earlier timing than the muon beam, where electrons contaminated the beam scattered at the target and surrounding materials. The time origin of the spectrum is defined as approximately 40 ns after the end timing of the second pulse. The spectrum was fitted by log-likelihood method using the following function consisting of three exponentials:
where is the lifetime of the natAg muonic atom, is the lifetime of the muonic atom of air and the acrylic case, is a long-lived background, and () are coefficients. The lifetimes of these three components were well separated and could be deduced by the free fitting. The second term originated from the events where the muon stopped in the air or the acrylic case, which constitutes approximately 2% of the total count. The air and acrylic contain carbon (C), nitrogen (N), and oxygen (O). Because their lifetimes are similar ( ns, ns, and ns [4]), the decay curve was represented with a single exponential with an averaged lifetime (). The value was determined to be 1.94(8) s. The quoted uncertainty contained the statistical uncertainty from the fitting (0.05 s) and the systematic uncertainty. The systematic uncertainty originates from variations in the ratio of muons stopped in the air and acrylic case, which depends on the shape of the target. Hence, an additional systematic uncertainty of 0.03 s was estimated in , which is used in subsequent analyses. The third term of Eq. (4) represents the long-lived background which was not related to the muon decay. Although the existence of such backgrounds has been known for a long time, its origin has remained unclear. It might be attributable to neutrons from the primary beam dump. The value was deduced from the fitting to be 25.8(16) s. The shape (decay constant) of the long-lived background was independent of the targets, so we used the same value for all targets. Since the events before the beam arrival were negligible, a constant term was not included in the fitting function. The lifetime of the muonic Ag was deduced to be 87.0(5) ns.
Figure 3 shows decay spectra measured with nine targets. The lifetimes of each target were deduced by fitting the following function:
where is the lifetime of the muonic atom of the target measured. Coefficients, (), are the parameters to be fitted with . If the target was a self-supported metal plate without using the Ag masks and backing plate, the decay component of Ag was omitted in the fitting, i.e., . For the self-supported metal targets, was deduced independently from the Ag sample measurement as discussed above. The primary background contribution, the second term of Eq. (4), was attributed to air, which contains less carbon. The lifetime of 1.88(5) s was used based on the composition of air and the muonic lifetime of nitrogen and oxygen. The muon stopping probability in the air was estimated to be approximately 1% using the GEANT4 toolkit [25, 26, 27]. This value was consistent with the stopping rate deduced from fits to data obtained with the self-supported metal targets. Note that the hyperfine splitting of muon state of 27Al can affect the lifetime of the muonic 27Al. Nevertheless, this effect was not considered; the decay curve was fitted with a single exponential function following the same analysis procedure as in Ref. [4].
The uncertainty in the lifetimes was estimated using the Monte Carlo method. In this method, the fixed parameters (, , and ) and fitting range were randomly varied within their respective uncertainties. The timing spectrum was then fit 500 times, resulting in a distribution of values. The standard deviation of this distribution was taken as the final uncertainty in . Table 1 summarizes the obtained lifetimes for each target.
IV Results
| Target | Time (h) | Huff factor | Lifetime (ns) | Capture rate (s-1) | ||||
| Nucleus | Form | Size (mm) | Weight (g) | Present | Suzuki et al.[4] | Present | ||
| 28Si | Powder in case | 15.02.8 | 0.500 | |||||
| 29Si | Powder in case | 15.02.8 | 0.500 | |||||
| 30Si | Small pieces in case | 15.02.8 | 0.500 | |||||
| natSi | Self-supported plate | 501.96 | 9.00 | |||||
| natSi | Powder in case | 15.02.8 | 0.550 | |||||
| natMg | Self-supported plate | 50502.0 | 8.54 | |||||
| 27Al | Self-supported plate | 50502.0 | 13.6 | |||||
| 27Al | Plate in case | 14.81.0 | 0.451 | |||||
| 55Mn | Chippings in case | 15.02.0 | 2.127 | |||||
| natAg | Plate in case | 14.80.3 | 0.544 | |||||
Table 1 summarized the measured lifetime and corresponding capture rates, calculated using Eq. (3). The Huff factors used in the conversion were taken from Ref. [4] and are included in the table. In this study, the lifetimes of isotopically enriched 28,29,30Si were obtained for the first time. The lifetimes of natMg, 27Al, 55Mn, and natAg were found to be consistent with previous measurement [4], confirming the reliability of our experimental setup and data analysis. However, the lifetime of natSi deviated from the previously reported value. To investigate the effect of the different forms of the targets and the acrylic case, we measured two natSi targets: a self-supported metal plate and a powder form encapsulated in the acrylic case. Both targets yielded consistent results, confirming that the shapes of the targets did not influence the measured lifetime. Additionally, the lifetime of natSi was also deduced from that of the enriched targets obtained in the present study. A Monte-Carlo calculation was used to create the decay curve of natSi by summing the timing spectrum of enriched Si, weighted by the natural abundance [28]. By fitting this simulated decay curve, the average lifetime of natSi was deduced to be 749.7(48) ns, consistent with our measurements as illustrated in Fig. 4. While our adopted value of 751.4(10) ns, an uncertainty-weighted average of the present results, deviates from the previous value of 756.0(10) ns, the reason for this discrepancy remains unclear.
V Discussions
| Target | Lifetime | Propagated uncertainty | Total | ||||
| Range | |||||||
| 28Si | 743.9 | 2.01 | 0.44 | 4.61 | 0.85 | 0.32 | 5.0 |
| 29Si | 794.2 | 1.85 | 0.39 | 2.89 | 0.91 | 0.35 | 3.5 |
| 30Si | 859.0 | 1.96 | 0.43 | 1.62 | 1.08 | 0.16 | 2.7 |
| natSi (plate) | 751.6 | 0.77 | - | 0.68 | 0.20 | 0.20 | 1.0 |
| natSi (in case) | 749.8 | 2.26 | 0.35 | 1.40 | 0.80 | 0.18 | 2.8 |
| natMg | 1065.5 | 1.96 | - | 0.39 | 0.27 | 0.36 | 2.0 |
| 27Al (plate) | 863.7 | 0.94 | - | 0.04 | 0.23 | 0.10 | 1.0 |
| 27Al (in case) | 862.5 | 3.19 | 0.43 | 7.58 | 0.98 | 0.34 | 8.1 |
| 55Mn | 230.3 | 3.57 | 0.66 | 2.54 | 0.42 | 0.96 | 4.5 |
In this study, we have obtained the muon nuclear capture rates of ten targets. As the present study was the first investigation of muonic lifetime using the SR spectrometer at MLF, we first discuss the measurement precision. Table 2 details the systematic breakdown of the uncertainty. Propagated uncertainties from each parameter were estimated using the Monte-Carlo method by keeping other parameters fixed. The statistical uncertainty on for most targets was approximately 1–2 ns. For targets in acrylic cases, the propagated uncertainty from dominated the total uncertainty. For the self-supported metal plate targets, namely natSi, natMg, and 27Al, the statistical uncertainty dominated the total uncertainty, with uncertainties propagated from and having a minor impact on the overall precision. The present results demonstrate the potential of our method to measure the lifetimes of various nuclei with a precision of 1–2 ns within a few hours of measurement.
Table 3 compares the capture rate obtained in the present study with theoretical calculations. The capture rates for natSi and 27Al in the table were calculated by taking an uncertainty-weighted average of two experimental data. The isotope dependence of the capture rate for the silicon isotopes is illustrated in Fig. 5. The data for natSi is also plotted at an average mass number of 28.11. The experimental capture rates decrease linearly with increasing mass numbers. This constant rate of decrease indicates that the capture rate is not influenced by the odd-even effect of the neutron number.
The isotope dependence of the capture rate was compared with the Primakoff [12] and Goulard-Primakoff [13] formulas, which are currently the only theoretical expressions available for comparison with experimental data in isotope dependence. The Primakoff formula is given by
where represents an overlap of the wave functions between the muon and nucleus, and and are the mass number and the atomic number of the target nucleus, respectively. While most previous studies and applications [4, 29, 30, 27, 14] have taken the values from original work by Ford and Wills [31], we use obtained by a theoretical calculation explained below. The parameters and are coefficients obtained from a global fit of the experimental data, where and taken from Ref. [4] were used in the following calculation. Equation (6) was later modified to introduce higher order corrections by Goulard and Primakoff [13] as
where the parameters , , , and were taken from TRIUMF data [4].
The capture rate is predominantly influenced by two effects: an overlap of the wave functions between the muon and the nucleus, and the Pauli exclusion effect, which is expressed as , and parameters of and in the Primakoff and Goulard-Primakoff formulas, respectively.
To investigate the isotope effect on , we performed a microscopic calculation using the definition proposed by Ford and Wills [31] as
where fm is the Bohr radius for muons [32], is the charge density distribution of the nucleus, and and are the radial component of the upper and lower components of the Dirac spinor of the muon in the orbital, respectively. Note that the normalization of the Dirac spinor is
The charge density distribution was calculated by using the Skyrme Hartree-Fock-Bogoliubov calculation [33, 34] with assuming the spherical symmetry, where of odd-mass nuclei was calculated using the equal filling approximation without considering the blocking or the time-odd component. The SGII effective interaction [35] and the volume-type pairing with the 60 MeV cutoff in the Hartree-Fock equivalent energy [36] were used for the particle-hole and particle-particle channel, respectively [37]. The Coulomb potential formed by , the vacuum polarization formed by the point charge [38, 39], and the reduced mass calculated based on the AME2020 [40, 41] were considered for and [15].
| Nucleus | Exp. | Primakoff | Others | ||
| 28Si | 0.968 [16] | ||||
| 0.823 [17] | |||||
| 0.789 [18] | |||||
| 0.892 [19] | |||||
| 29Si | |||||
| 30Si | |||||
| natSi | 111Average of two experimental data in Table 1. | 222Weighted average of by the natual abundance of each isotopes [28]. | |||
| natMg | 222Weighted average of by the natual abundance of each isotopes [28]. | 0.612 [16] | |||
| 0.454 [17] | |||||
| 27Al | 111Average of two experimental data in Table 1. | 0.698 [16] | |||
| 55Mn | |||||
| natAg | 222Weighted average of by the natual abundance of each isotopes [28]. | ||||
Table 3 lists the calculated values. For the silicon isotopes, the calculated is slightly larger than the original value of [31]. Because this constant value was used to deduce parameters of and in the Primakoff and Goulard-Primakoff formulas [4], respectively, absolute values of the calculated capture rates by these formulas may be changed by reevaluating the parameters using our values. Thus, the isotope dependence of the capture rate is discussed with relative values normalized to that of 28Si. While the reductions of are 9.54(12)% and 20.18(23)% for 29Si and 30Si, respectively, the corresponding reductions in are two order of magnitude smaller of 0.10% and 0.22%. Thus, the effect of on the isotope dependence of is negligible for the Si isotopes, in contrast to an Eu isotope case [8]. The significant isotope effect from between 151Eu88 and 153Eu90 might be due to the large change in nuclear deformation in the region of the neutron number from 82 to 92. In the case of the Si isotopes, the quadrupole deformation parameters () were experimentally deduced from the transition probability between and to be 0.407(7) and 0.315(7) for 28Si and 30Si, respectively [42]. The nuclear deformation affects through a root-mean-squere charge radius () of through Eq. (8). The experimenatal were 3.1224(24) and 3.1336(40) fm in Ref. [43] and our calculated of were 3.15 and 3.16 fm for 28Si and 30Si, respectively. The calculated is consistent with the experimental values, although the calculation assumed spherical symmetry. Therefore, we conclude that the isotope dependence of for the Si isotopes is not explained by that of .
The Pauli exclusion effect was expressed in the term with the parameter and in the Primakoff and Goulard-Primakoff formulas, respectively. While the absolute capture rates calculated by the Primakoff formula underpredict experimental data, the isotope dependence is reasonably well reproduced, as shown in Fig. 5. The Goulard-Primakoff formula overestimates the capture rates of 28Si and natSi, while it underestimates those of 29Si and 30Si. The inadequacy of both Primakoff and Goulard-Primakoff formulas in accounting for the isotope effect has also been previously pointed out in Refs. [4, 14]. A re-evaluation of these Primakoff’s formulas is a worthwhile direction for future research, and the experimental method established in this study offers an opportunity for systematic measurements.
Capture rates obtained by other theories [16, 17, 18, 19] are also listed in Table 3. The global systematics of the experimental capture rates and comparisons with those theories are out of the scope of the present study and are discussed in a separated publication [14].
The measured isotope dependence of the muon capture rate was also compared with a microscopic and evaporation model (MEM) calculation [15]. In this model, the wave function of the muonic atom was calculated using the same method as that used to calculate . The muon capture rates were calculated using the second Tamm-Dancoff approximation [44] with effective interactions of SkO’ [45] and SGII [35]. Calculated capture rates strongly depend on the axial vector and induced pseudo-scalar coupling constants ( and , respectively), as well as the choice of effective interaction. As the capture rate calculated by MEM has large ambiguity due to the uncertainties in and in medium, possible combinations of the parameter set of and were deduced to reproduce the measured capture rates. Note that the values of and presented in this study are at zero momentum transfer, namely and , respectively. Due to the limitation of the MEM calculation to even-even nuclei, we discuss 28Si and 30Si hereinafter.
Figure 6 shows a variation of the muon capture rates by changing and using the SKO’ and SGII effective interactions. The solid lines represent the calculated capture rates for a given value as functions of . The range of was set to 1.0–1.3 and 0.7–1.3 for SkO’ and SGII, respectively, which include the value of free nucleon () and typical quenched value () [46]. The range of was set to cover the range to reproduce the experimental result with the given range, including the free nucleon value with the muon capture on the hydrogen of [47] and other previous investigations [48, 1, 49]. The muon capture rates are raised with increasing and decrease with increasing in all cases. The experimental results of capture rates are also shown in Fig. 6 with the red lines. By taking an overlap between the experimental values and calculated ones, parameter sets of and , which reproduce the experimental muon capture rates, were obtained, as illustrated in Fig. 7. In the figure, the lines represent the (, ) set constrained by experimental results, and the 1 uncertainties of the experimental values are shown with colored shade. The Goldberger-Treiman relation, which expresses the relation between and based on the assumption of Partially Conserved Axial Current (PCAC) and the pion-pole dominance of the induced pseudoscalar form factor, yields [50], as shown by the black solid line in the figure.
The values smaller than that of free nucleon of 1.27 are required to reproduce the experimental results with the given range of , suggesting the quenching of . When comparing the parameter set between 28Si and 30Si with the same effective interactions, there were no overlapping of parameter sets, indicating the isotope dependence of quenching on both or either of and . The larger quenching of was required for 30Si than 28Si when was fixed, consistent with the trend discussed in Refs. [51, 52, 53, 46]. A comparison of constrained (, ) regions for two effective interactions, SkO’ and SGII, revealed that the calculation using SkO’ reproduced the capture rate with smaller quenching than that using SGII. This discrepancy can be understood by different energy distributions populated by muon capture. The excited states calculated with SkO’ distribute at higher energies compared to SGII, and phase spaces available for muon neutrinos in the final state are limited. This results in the smaller muon capture rate of SkO’ than SGII [44], leading to the smaller quenching.
VI Conclusion
We have measured the lifetimes of the muonic atom using a SR spectrometer at the muon facility in MLF, J-PARC. In this study, the lifetimes of isotopically enriched 28,29,30Si were obtained for the first time. The lifetimes of natMg, 27Al, 55Mn, and natAg were also measured and found to be consistent with previous measurements, confirming the reliability of the experimental method and analysis procedure.
We compared the isotope dependence of the lifetimes for 28,29,30Si with the Primakoff and Goulard-Primakoff formulas, utilizing newly calculated values. The isotope dependence of is negligibly small and does not explain that of the capture rate. While the absolute lifetimes calculated by the Primakoff formula exceed our experimental data, the isotope dependence is reasonably well-reproduced. The Goulard-Primakoff formula accurately predicts the lifetimes of 28Si and natSi but fails to reproduce those of 29Si and 30Si.
Through a comparison of our experimental results on 28,30Si with the recent microscopic and evaporation model (MEM) using two effective interactions (SkO’ and SGII), we constrained the axial vector () and induced pseudoscalar () coupling constants of the weak interaction. The constrained values of are found to be smaller than the free nucleon value of 1.27, suggesting the quenching of in the MEM calculation. The amount of quenching exhibits isotope dependence between 28Si and 30Si. Furthermore, the choice of effective interaction also influences the required quenching of and .
The present experiment establishes a method for measuring the lifetimes of the muonic atom at MLF in J-PARC, opening up a new possibility for systematic investigations of capture rates. The re-evaluation of the muon capture rates with the Primakoff and Goulard-Primakoff formulas will provide better estimation in general-purpose simulation codes such as GEANT4 [27], PHITS [29], and FLUKA [30]. The systematic study of the capture rate will contribute to a further understanding of the quenching of and of the nuclear medium.
Acknowledgements.
We are grateful for the valuable discussions with Dr. T. Y. Saito (RIKEN), Dr. H. Iwamoto (JAEA), and Dr. T. Miyagi (Univ. Tsukuba). The muon experiment at the Materials and Life Science Experimental Facility of the J-PARC was performed under a user program (Proposal No. 2024MI01). The present work is partially supported by the JSPS Grant-in-Aid under Grant Nos. JP19H05664, JP22K20372, JP23H04526, JP23H01845, JP23K01845, JP23K03426, JP24H00073, JP24K00647, and JP24K17057. R.M. is supported by the Force-front Physics and Mathematics Program in Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) program, and the JSR Fellowship from the University of Tokyo. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program. The numerical calculations were partially performed on cluster computers at the RIKEN iTHEMS program.References
- Measday [2001] D. F. Measday, Phys. Rep. 354, 243 (2001).
- Navas et al. [2024] S. Navas, C. Amsler, T. Gutsche, C. Hanhart, J. J. Hernández-Rey, C. Lourenço, A. Masoni, M. Mikhasenko, R. E. Mitchell, C. Patrignani, C. Schwanda, S. Spanier, G. Venanzoni, C. Z. Yuan, K. Agashe, G. Aielli, B. C. Allanach, J. Alvarez-Muñiz, M. Antonelli, E. C. Aschenauer, D. M. Asner, K. Assamagan, H. Baer, S. Banerjee, R. M. Barnett, L. Baudis, C. W. Bauer, J. J. Beatty, J. Beringer, A. Bettini, O. Biebel, K. M. Black, E. Blucher, R. Bonventre, R. A. Briere, A. Buckley, V. D. Burkert, M. A. Bychkov, R. N. Cahn, Z. Cao, M. Carena, G. Casarosa, A. Ceccucci, A. Cerri, R. S. Chivukula, G. Cowan, K. Cranmer, V. Crede, O. Cremonesi, G. D’Ambrosio, T. Damour, D. de Florian, A. de Gouvêa, T. DeGrand, S. Demers, Z. Demiragli, B. A. Dobrescu, M. D’Onofrio, M. Doser, H. K. Dreiner, P. Eerola, U. Egede, S. Eidelman, A. X. El-Khadra, J. Ellis, S. C. Eno, J. Erler, V. V. Ezhela, A. Fava, W. Fetscher, B. D. Fields, A. Freitas, H. Gallagher, T. Gershon, Y. Gershtein, T. Gherghetta, M. C. Gonzalez-Garcia, M. Goodman, C. Grab, A. V. Gritsan, C. Grojean, D. E. Groom, M. Grünewald, A. Gurtu, H. E. Haber, M. Hamel, S. Hashimoto, Y. Hayato, A. Hebecker, S. Heinemeyer, K. Hikasa, J. Hisano, A. Höcker, J. Holder, L. Hsu, J. Huston, T. Hyodo, A. Ianni, M. Kado, M. Karliner, U. F. Katz, M. Kenzie, V. A. Khoze, S. R. Klein, F. Krauss, M. Kreps, P. Križan, B. Krusche, Y. Kwon, O. Lahav, L. P. Lellouch, J. Lesgourgues, A. R. Liddle, Z. Ligeti, C.-J. Lin, C. Lippmann, T. M. Liss, A. Lister, L. Littenberg, K. S. Lugovsky, S. B. Lugovsky, A. Lusiani, Y. Makida, F. Maltoni, A. V. Manohar, W. J. Marciano, J. Matthews, U.-G. Meißner, I.-A. Melzer-Pellmann, P. Mertsch, D. J. Miller, D. Milstead, K. Mönig, P. Molaro, F. Moortgat, M. Moskovic, N. Nagata, K. Nakamura, M. Narain, P. Nason, A. Nelles, M. Neubert, Y. Nir, H. B. O’Connell, C. A. J. O’Hare, K. A. Olive, J. A. Peacock, E. Pianori, A. Pich, A. Piepke, F. Pietropaolo, A. Pomarol, S. Pordes, S. Profumo, A. Quadt, K. Rabbertz, J. Rademacker, G. Raffelt, M. Ramsey-Musolf, P. Richardson, A. Ringwald, D. J. Robinson, S. Roesler, S. Rolli, A. Romaniouk, L. J. Rosenberg, J. L. Rosner, G. Rybka, M. G. Ryskin, R. A. Ryutin, B. Safdi, Y. Sakai, S. Sarkar, F. Sauli, O. Schneider, S. Schönert, K. Scholberg, A. J. Schwartz, J. Schwiening, D. Scott, F. Sefkow, U. Seljak, V. Sharma, S. R. Sharpe, V. Shiltsev, G. Signorelli, M. Silari, F. Simon, T. Sjöstrand, P. Skands, T. Skwarnicki, G. F. Smoot, A. Soffer, M. S. Sozzi, C. Spiering, A. Stahl, Y. Sumino, F. Takahashi, M. Tanabashi, J. Tanaka, M. Taševský, K. Terao, K. Terashi, J. Terning, U. Thoma, R. S. Thorne, L. Tiator, M. Titov, D. R. Tovey, K. Trabelsi, P. Urquijo, G. Valencia, R. Van de Water, N. Varelas, L. Verde, I. Vivarelli, P. Vogel, W. Vogelsang, V. Vorobyev, S. P. Wakely, W. Walkowiak, C. W. Walter, D. Wands, D. H. Weinberg, E. J. Weinberg, N. Wermes, M. White, L. R. Wiencke, S. Willocq, C. L. Woody, R. L. Workman, W.-M. Yao, M. Yokoyama, R. Yoshida, G. Zanderighi, G. P. Zeller, R.-Y. Zhu, S.-L. Zhu, F. Zimmermann, P. A. Zyla, J. Anderson, M. Kramer, P. Schaffner, W. Zheng, and Particle Data Group Collaboration, Phys. Rev. D. 110, 030001 (2024).
- Huff [1961] R. W. Huff, Ann. Phys. 16, 288 (1961).
- Suzuki et al. [1987] T. Suzuki, D. F. Measday, and J. P. Roalsvig, Phys. Rev. C. 35, 2212 (1987).
- Bertram et al. [1960] W. J. Bertram, R. A. Reiter, T. A. Romanowski, and R. B. Sutton, Phys. Rev. Lett. 5, 61 (1960).
- Cramer et al. [1962] W. A. Cramer, V. L. Telegdi, R. Winston, and R. A. Lundy, Il Nuovo Cimento 24, 546 (1962).
- Povel [1970] H. Povel, Phys. Lett. B 33, 620 (1970).
- Petitjean et al. [1971] C. Petitjean, H. Backe, R. Engfer, U. Jahnke, K. H. Lindenberger, H. Schneuwly, W. U. Schröder, and H. K. Walter, Nucl. Phys. A 178, 193 (1971).
- Bobrov et al. [1966] V. Bobrov, V. Varlamov, Y. Grashin, B. Dolgoshein, V. Kirillov-Ugryumov, Y. Nikitin, V. Roganov, A. Samoilov, and S. Somov, Sov. J. Nucl. Phys. 4, 53 (1966).
- Mamedov et al. [2000] T. N. Mamedov, V. G. Grebinnik, K. I. Gritsai, V. N. Duginov, V. A. Zhukov, V. G. Ol’shevskii, and A. V. Stoikov, JETP Lett. 71, 451 (2000).
- Fynbo et al. [2003] H. O. U. Fynbo, V. Egorov, V. Brudanin, M. Chirtchenko, J. Deutsch, V. Lebedev, C. Petitjean, K. Riisager, and S. Vassiliev, Nucl. Phys. A 724, 493 (2003).
- Primakoff [1959] H. Primakoff, Rev. Mod. Phys. 31, 802 (1959).
- Goulard and Primakoff [1974] B. Goulard and H. Primakoff, Phys. Rev. C 10, 2034 (1974).
- Iwamoto et al. [2025] H. Iwamoto, M. Niikura, and R. Mizuno, Phys. Rev. C. 111, 034614 (2025).
- Minato et al. [2023] F. Minato, T. Naito, and O. Iwamoto, Phys. Rev. C. 107, 054314 (2023).
- Chiang et al. [1990] H. C. Chiang, E. Oset, and P. F. De Córdoba, Nucl. Phys. A 510, 591 (1990).
- Zinner et al. [2006] N. T. Zinner, K. Langanke, and P. Vogel, Phys. Rev. C. 74, 024326 (2006).
- Marketin et al. [2009] T. Marketin, N. Paar, T. Nikšić, and D. Vretenar, Phys. Rev. C 79, 054323 (2009).
- Giannaka and Kosmas [2015] P. G. Giannaka and T. S. Kosmas, Phys. Rev. C 92, 014606 (2015).
- Mizuno et al. [2025] R. Mizuno, M. Niikura, T. Y. Saito, T. Matsuzaki, S. Abe, H. Fukuda, M. Hashimoto, A. Hillier, K. Ishida, N. Kawamura, S. Kawase, T. Kawata, K. Kitafuji, F. Minato, M. Oishi, A. Sato, K. Shimomura, P. Strasser, S. Takeshita, D. Tomono, and Y. Watanabe, arXiv:2507.19753 [nucl-ex] (2025).
- Miyake et al. [2012] Y. Miyake, K. Shimomura, N. Kawamura, P. Strasser, A. Koda, S. Makimura, H. Fujimori, Y. Ikedo, K. Nakahara, S. Takeshita, M. Kato, K. Kojima, Y. Kobayashi, K. Nishiyama, R. Kadono, W. Higemoto, T. U. Ito, K. Ninomiya, K. Kubo, and K. Nagamine, Phys. Procedia 30, 46 (2012).
- Kojima et al. [2014] K. M. Kojima, T. Murakami, Y. Takahashi, H. Lee, S. Y. Suzuki, A. Koda, I. Yamauchi, M. Miyazaki, M. Hiraishi, H. Okabe, S. Takeshita, R. Kadono, T. Ito, W. Higemoto, S. Kanda, Y. Fukao, N. Saito, M. Saito, M. Ikeno, T. Uchida, and M. M. Tanaka, J. Phys. Conf. Ser. 551, 012063 (2014).
- Wakata et al. [2025] M. Wakata, S. Akamatsu, T. Fujiie, T. Furuyama, L. Hara, Y. Ishikawa, T. Ito, T. Kikuchi, T. Mibe, S. Ozaki, M. Yokomizo, and J. Murata, Prog. Theor. Exp. Phys. 2025, 023H01 (2025).
- Takeshita et al. [2024] S. Takeshita, I. Umegaki, M. Tampo, P. Strasser, Y. Ikedo, A. Koda, W. Higemoto, T. Yuasa, N. Kawamura, Y. Miyake, and K. Shimomura, Interactions 245, 1 (2024).
- Agostinelli et al. [2003] S. Agostinelli, J. Allison, K. Amako, J. Apostolakis, H. Araujo, P. Arce, M. Asai, D. Axen, S. Banerjee, G. Barrand, F. Behner, L. Bellagamba, J. Boudreau, L. Broglia, A. Brunengo, H. Burkhardt, S. Chauvie, J. Chuma, R. Chytracek, G. Cooperman, G. Cosmo, P. Degtyarenko, A. Dell’Acqua, G. Depaola, D. Dietrich, R. Enami, A. Feliciello, C. Ferguson, H. Fesefeldt, G. Folger, F. Foppiano, A. Forti, S. Garelli, S. Giani, R. Giannitrapani, D. Gibin, J. J. Gómez Cadenas, I. González, G. Gracia Abril, G. Greeniaus, W. Greiner, V. Grichine, A. Grossheim, S. Guatelli, P. Gumplinger, R. Hamatsu, K. Hashimoto, H. Hasui, A. Heikkinen, A. Howard, V. Ivanchenko, A. Johnson, F. W. Jones, J. Kallenbach, N. Kanaya, M. Kawabata, Y. Kawabata, M. Kawaguti, S. Kelner, P. Kent, A. Kimura, T. Kodama, R. Kokoulin, M. Kossov, H. Kurashige, E. Lamanna, T. Lampén, V. Lara, V. Lefebure, F. Lei, M. Liendl, W. Lockman, F. Longo, S. Magni, M. Maire, E. Medernach, K. Minamimoto, P. Mora de Freitas, Y. Morita, K. Murakami, M. Nagamatu, R. Nartallo, P. Nieminen, T. Nishimura, K. Ohtsubo, M. Okamura, S. O’Neale, Y. Oohata, K. Paech, J. Perl, A. Pfeiffer, M. G. Pia, F. Ranjard, A. Rybin, S. Sadilov, E. Di Salvo, G. Santin, T. Sasaki, N. Savvas, Y. Sawada, S. Scherer, S. Sei, V. Sirotenko, D. Smith, N. Starkov, H. Stoecker, J. Sulkimo, M. Takahata, S. Tanaka, E. Tcherniaev, E. Safai Tehrani, M. Tropeano, P. Truscott, H. Uno, L. Urban, P. Urban, M. Verderi, A. Walkden, W. Wander, H. Weber, J. P. Wellisch, T. Wenaus, D. C. Williams, D. Wright, T. Yamada, H. Yoshida, and D. Zschiesche, Nucl. Instrum. Methods Phys. Res. A 506, 250 (2003).
- Allison et al. [2006] J. Allison, K. Amako, J. Apostolakis, H. Araujo, P. Arce Dubois, M. Asai, G. Barrand, R. Capra, S. Chauvie, R. Chytracek, G. A. P. Cirrone, G. Cooperman, G. Cosmo, G. Cuttone, G. G. Daquino, M. Donszelmann, M. Dressel, G. Folger, F. Foppiano, J. Generowicz, V. Grichine, S. Guatelli, P. Gumplinger, A. Heikkinen, I. Hrivnacova, A. Howard, S. Incerti, V. Ivanchenko, T. Johnson, F. Jones, T. Koi, R. Kokoulin, M. Kossov, H. Kurashige, V. Lara, S. Larsson, F. Lei, O. Link, F. Longo, M. Maire, A. Mantero, B. Mascialino, I. McLaren, P. Mendez Lorenzo, K. Minamimoto, K. Murakami, P. Nieminen, L. Pandola, S. Parlati, L. Peralta, J. Perl, A. Pfeiffer, M. G. Pia, A. Ribon, P. Rodrigues, G. Russo, S. Sadilov, G. Santin, T. Sasaki, D. Smith, N. Starkov, S. Tanaka, E. Tcherniaev, B. Tome, A. Trindade, P. Truscott, L. Urban, M. Verderi, A. Walkden, J. P. Wellisch, D. C. Williams, D. Wright, and H. Yoshida, IEEE Trans. Nucl. Sci. 53, 270 (2006).
- Allison et al. [2016] J. Allison, K. Amako, J. Apostolakis, P. Arce, M. Asai, T. Aso, E. Bagli, A. Bagulya, S. Banerjee, G. Barrand, B. R. Beck, A. G. Bogdanov, D. Brandt, J. M. C. Brown, H. Burkhardt, P. Canal, D. Cano-Ott, S. Chauvie, K. Cho, G. A. P. Cirrone, G. Cooperman, M. A. Cortés-Giraldo, G. Cosmo, G. Cuttone, G. Depaola, L. Desorgher, X. Dong, A. Dotti, V. D. Elvira, G. Folger, Z. Francis, A. Galoyan, L. Garnier, M. Gayer, K. L. Genser, V. M. Grichine, S. Guatelli, P. Guèye, P. Gumplinger, A. S. Howard, I. Hřivnáčová, S. Hwang, S. Incerti, A. Ivanchenko, V. N. Ivanchenko, F. W. Jones, S. Y. Jun, P. Kaitaniemi, N. Karakatsanis, M. Karamitros, M. Kelsey, A. Kimura, T. Koi, H. Kurashige, A. Lechner, S. B. Lee, F. Longo, M. Maire, D. Mancusi, A. Mantero, E. Mendoza, B. Morgan, K. Murakami, T. Nikitina, L. Pandola, P. Paprocki, J. Perl, I. Petrović, M. G. Pia, W. Pokorski, J. M. Quesada, M. Raine, M. A. Reis, A. Ribon, A. Ristić Fira, F. Romano, G. Russo, G. Santin, T. Sasaki, D. Sawkey, J. I. Shin, I. I. Strakovsky, A. Taborda, S. Tanaka, B. Tomé, T. Toshito, H. N. Tran, P. R. Truscott, L. Urban, V. Uzhinsky, J. M. Verbeke, M. Verderi, B. L. Wendt, H. Wenzel, D. H. Wright, D. M. Wright, T. Yamashita, J. Yarba, and H. Yoshida, Nucl. Instrum. Methods Phys. Res. A 835, 186 (2016).
- Meija et al. [2016] J. Meija, T. B. Coplen, M. Berglund, W. A. Brand, P. De Bièvre, M. Gröning, N. E. Holden, J. Irrgeher, R. D. Loss, T. Walczyk, and T. Prohaska, Pure Appl. Chem. 88, 293 (2016).
- Abe and Sato [2017] S. Abe and T. Sato, J. Nucl. Sci. Technol. 54, 101 (2017).
- Battistoni et al. [2015] G. Battistoni, T. Boehlen, F. Cerutti, P. W. Chin, L. S. Esposito, A. Fassò, A. Ferrari, A. Lechner, A. Empl, A. Mairani, A. Mereghetti, P. G. Ortega, J. Ranft, S. Roesler, P. R. Sala, V. Vlachoudis, and G. Smirnov, Ann. Nucl. Energy 82, 10 (2015).
- Ford and Wills [1962] K. W. Ford and J. G. Wills, Nucl. Phys. 35, 295 (1962).
- Tiesinga et al. [2021] E. Tiesinga, P. J. Mohr, D. B. Newell, and B. N. Taylor, Rev. Mod. Phys. 93, 025010 (2021).
- Vautherin and Brink [1972] D. Vautherin and D. M. Brink, Phys. Rev. C 5, 626 (1972).
- Dobaczewski et al. [1984] J. Dobaczewski, H. Flocard, and J. Treiner, Nucl. Phys. A 422, 103 (1984).
- Van Giai and Sagawa [1981] N. Van Giai and H. Sagawa, Phys. Lett. B 106, 379 (1981).
- Stoitsov et al. [2013] M. V. Stoitsov, N. Schunck, M. Kortelainen, N. Michel, H. Nam, E. Olsen, J. Sarich, and S. Wild, Comput. Phys. Commun. 184, 1592 (2013).
- Naito et al. [2023] T. Naito, T. Oishi, H. Sagawa, and Z. Wang, Phys. Rev. C 107, 054307 (2023).
- Uehling [1935] E. A. Uehling, Phys. Rev. 48, 55 (1935).
- Wayne Fullerton and Rinker [1976] L. Wayne Fullerton and G. A. Rinker, Phys. Rev. A 13, 1283 (1976).
- Huang et al. [2021] W. J. Huang, M. Wang, F. G. Kondev, G. Audi, and S. Naimi, Chin. Phys. C 45, 030002 (2021).
- Wang et al. [2021] M. Wang, W. J. Huang, F. G. Kondev, G. Audi, and S. Naimi, Chin. Phys. C 45, 030003 (2021).
- Raman et al. [2001] S. Raman, C. W. Nestor, Jr, and P. Tikkanen, At. Data Nucl. Data Tables 78, 1 (2001).
- Angeli and Marinova [2013] I. Angeli and K. P. Marinova, At. Data Nucl. Data Tables 99, 69 (2013).
- Minato [2016] F. Minato, Phys. Rev. C. 93, 044319 (2016).
- Reinhard et al. [1999] P.-G. Reinhard, D. J. Dean, W. Nazarewicz, J. Dobaczewski, J. A. Maruhn, and M. R. Strayer, Phys. Rev. C. 60, 014316 (1999).
- Suhonen [2017] J. T. Suhonen, Frontiers in Physics 5, 55 (2017), and references theirin.
- Andreev et al. [2013] V. A. Andreev, T. I. Banks, R. M. Carey, T. A. Case, S. M. Clayton, K. M. Crowe, J. Deutsch, J. Egger, S. J. Freedman, V. A. Ganzha, T. Gorringe, F. E. Gray, D. W. Hertzog, M. Hildebrandt, P. Kammel, B. Kiburg, S. Knaack, P. A. Kravtsov, A. G. Krivshich, B. Lauss, K. R. Lynch, E. M. Maev, O. E. Maev, F. Mulhauser, C. Petitjean, G. E. Petrov, R. Prieels, G. N. Schapkin, G. G. Semenchuk, M. A. Soroka, V. Tishchenko, A. A. Vasilyev, A. A. Vorobyov, M. E. Vznuzdaev, P. Winter, and MuCap Collaboration, Phys. Rev. Lett. 110, 012504 (2013).
- Jokiniemi et al. [2019] L. Jokiniemi, J. Suhonen, H. Ejiri, and I. Hashim, Physics Letters B 794, 143 (2019).
- Gorringe and Fearing [2003] T. Gorringe and H. W. Fearing, Rev. Mod. Phys. 76, 31 (2003).
- Goldberger and Treiman [1958] M. L. Goldberger and S. B. Treiman, Phys. Rev. 111, 354 (1958).
- Fearing and Welsh [1992] H. W. Fearing and M. S. Welsh, Phys. Rev. C 46, 2077 (1992).
- Gorringe et al. [1998] T. P. Gorringe, D. S. Armstrong, C. Q. Chen, E. Christy, B. C. Doyle, P. Gumplinger, H. W. Fearing, M. D. Hasinoff, M. A. Kovash, and D. H. Wright, Phys. Rev. C 58, 1767 (1998).
- Barea et al. [2013] J. Barea, J. Kotila, and F. Iachello, Phys. Rev. C 87, 014315 (2013).